OpenStudy (anonymous):

Find the quotient. Simplify if possible. 7over9÷ 5over6 =

OpenStudy (anonymous):

again... dividing by a fraction is the same as multiplying by its reciprocal

OpenStudy (anonymous):

ex. \[\frac{1}{2} \div \frac{3}{4} = \frac{1}{2} \times \frac{ 4}{3}\]

OpenStudy (anonymous):

if you're having issues multiplying fractions ex. \[\frac{1}{2} \times \frac{5}{4} = \frac { 1 \times 5}{2 \times 4}\]

OpenStudy (anonymous):

5